### v3

# Yuang Li, Yinglu Li, Min Zhang, Chang Su, Mengxin Ren, Xiaosong Qiao, Xiaofeng Zhao, Mengyao Piao, Jiawei Yu, Xinglin Lv, Miaomiao Ma, Yanqing Zhao, Hao Yang

Abstract
End-to-end automatic speech recognition (ASR) systems often struggle to recognize rare name entities, such as personal names, organizations, and terminologies not frequently encountered in the training data. This paper presents Contextual Biasing Whisper (CB-Whisper), a novel ASR system based on OpenAI's Whisper model that can recognize user-defined name entities by performing open-vocabulary keyword-spotting (OV-KWS) using the hidden states of Whisper encoder. The recognized entities are used as prompts for the Whisper decoder. We first propose a multitask training approach with OV-KWS and ASR tasks to optimize the model. Experiments show that this approach substantially improves the entity recalls compared to the original Whisper model on Chinese Aishell hot word subsets and two internal code-switch test sets. However, we observed a slight increase in mixed-error-rate (MER) on internal test sets due to catastrophic forgetting. To address this problem and use different sizes of the Whisper model without finetuning, we propose to use OV-KWS as a separate module and construct a spoken form prompt to prevent hallucination. The OV-KWS module consistently improves MER and Entity Recall for whisper-small, medium, and large models.
Index Terms: speech recognition, contextual biasing, keyword-spotting
1 Introduction
End-to-end (E2E) automatic speech recognition (ASR) models [1, 2, 3] have gained popularity in recent years for their simplicity and unified architecture. However, they suffer from low recall of proper nouns that are rare in the training data. Contextual biasing (CB) is a technique that leverages external knowledge of the expected words in the speech input to mitigate the long-tail word problem. One of the early approaches is shallow fusion [4, 5], which builds an n-gram finite state transducer from a predefined list of hot words and boosts their scores during beam search decoding. However, this approach has some drawbacks, such as the challenge of finding the optimal fusion weight. Therefore, researchers have proposed deep fusion methods [6, 7, 8, 9, 10], which train a contextual encoder together with the ASR model from scratch. These methods embed the entity words and the speech signal into the same feature space, and the decoder uses both contextual and acoustic information to generate the transcription. To adapt existing ASR models, some methods use adaptors [11, 12] to modify the intermediate features of the ASR model or pointer networks [13] to modify the output distributions. Moreover, shallow and deep fusion methods can be combined to further enhance the performance [14, 15].
OpenAI's Whisper [16] is a state-of-the-art ASR model based on the Transformer [17] architecture and it was trained on a large-scale dataset of 680k hours of speech data. Unlike previous models that require a separate biasing module, it can adaptively bias the generation of entity words by providing a simple prompt to the decoder. In this paper, we propose a novel method called contextual biasing Whisper (CB-Whisper) which incorporates an open-vocabulary keyword-spotting (OV-KWS) module between the encoder and the decoder of the Whisper to search entity words before constructing prompts. The OV-KWS module is inspired by vision-based KWS methods [18, 19] and uses features from the Whisper encoder to construct a cosine similarity matrix. To train the OV-KWS model and enhance Whisper's ability to understand prompts, we introduce a multitask training approach with OV-KWS and ASR tasks. Our system enhances entity recalls with absolute improvements of up to 80% on Aishell hot word subsets [20] and up to 10% on Internal code-switch datasets. However, we noticed that CB-Whisper finetuned on a small-scale dataset has higher mixed-error-rates (MERs) compared to the Whisper model on real-world Internal datasets due to catastrophic forgetting. Therefore, we propose to freeze the Whisper model and use OV-KWS as an independent module. To align the format of the prompt with the original Whisper training setup, we devised a spoken form prompt that mimics the history transcription. For whisper-small, medium, and large models, our system can achieve absolute MER reductions of 2.4%, 1.4%, and 1.8% respectively on Internal datasets. Note that the OV-KWS is a lightweight module using the Whisper-small encoder, making it computationally efficient.
Figure 1: (a) Entity Features are extracted using TTS followed by Whisper encoder. (b) The flowchart of the CB-Whisper model. (c) Illustration of OV-KWS module.
2 Methodology
2.1 System overview
Before the system can be deployed, a database is created by extracting acoustic representations for pre-defined entity words. As shown in Figure 1 (a), the speech signal for each entity word is generated with a Text-to-Speech (TTS) model. Then, the whisper encoder is used to obtain hidden states that will be later matched with the features of the input utterance.
As illustrated in Figure 1 (b), the proposed CB-Whisper identifies entity words before the decoder by utilizing the hidden states of the encoder. In detail, we first obtain the hidden representations of the input utterance by summing up the hidden states from all Whisper encoder layers with a learnable weight for each layer. Next, we calculate the cosine similarity matrix between the input utterance features and the stored features of each entity word. A binary Convolutional Neural Network (CNN) classifier then determines if the entity word is present in the utterance by recognizing the diagonal pattern in the similarity matrix (Figure 1 (c)). Finally, the predicted entity words are used as prompts to guide the whisper decoder so that these words can be recognized more accurately.
2.2 Open-vocabulary keyword-spotting
In the proposed system, OV-KWS is formulated as a binary classification task. A deep neural network predicts the presence of an entity word in the utterance by taking the cosine similarity matrix between the hidden states of entity words and the input utterance as input. As shown in Figure 1 (c), if an entity word exists in the utterance, a diagonal pattern can be observed. Since such a pattern is a local feature, we choose CNN as the classification model. The usage of the similarity matrix rather than the original hidden states improves the model's generalisability and allows the detection of words that were not seen during training. Following the common approach to using multi-level self-supervised features for downstream tasks [21], we use the weighted sum of multi-layer features from Whisper-encoder instead of a single-layer feature.
The OV-KWS module has a linear computational complexity to the number of keywords. To enhance the efficiency, we employed a very lightweight CNN network with only four layers and 0.2 million parameters. Furthermore, multiple keywords can be batched and executed in parallel on GPU.
An easy way to construct the dataset for the CNN classifier is to select positive words according to the transcription and to choose negative words randomly. However, this approach is too simplistic, as the negative samples may have significant differences from the positive samples, resulting in overfitting. Thus, we adopted hard negative samples [22], which are more likely to have overlaps with positive samples, increasing the difficulty of the classification task.
2.3 Multitask finetuning
The CNN classifier's capacity is limited by its small size. Therefore, it is necessary to finetune the Whisper model for the OV-KWS task while preserving its ASR ability. Hence, we propose to finetune all parameters of the Whisper model on the OV-KWS task and the ASR task jointly with the loss function shown in Equation 1. Additionally, during finetuning, we incorporate words from the ground-truth transcription into the prompt following the naive prompt format in Table 1. This enables the Whisper decoder to better understand the prompt format and leverage the word in the prompt.
$Loss=\mathcal{L}_{ASR}+\alpha\times\mathcal{L}_{KWS}$ (1)
where the loss is the weighted sum of ASR loss $\mathcal{L}_{ASR}$ and KWS classification loss $\mathcal{L}_{KWS}$. The weight $\alpha$ was set to three during training.
2.4 Prompting Whisper decoder
Prompt is a term that refers to a segment of text that provides the context, or the objective for the generation of large language models. Whisper, an encoder-decoder model, also supports prompting because the decoder can be viewed as an internal language model. The decoder used the transcription of the previous utterance as a prompt during training to incorporate contextual information for the recognition of the following utterance. During inference, any text-only prompt that is relevant to the input utterance can serve as historical context for decoding.
Table 1: Prompts for whisper decoder.
naive prompt
实体1, 实体2, 实体3
entity 1, entity 2, entity 3
spoken form prompt
今天演讲的主题是这个呃,实体1、实体2、实体3。好,那我就继续讲。
The topic of today’s speech is, ah, entity 1, entity 2, entity 3. Okay, then I’ll continue.
For the multitask training experiments, we adopt the naive prompt format, which consists of the plain concatenation of all the entity words. This format works well for the in-domain test sets (Aishell), but it may induce hallucination on more difficult spontaneous talks that involve code-switching situations. Moreover, the finetuned model’s ASR performance may slightly degrade, as it was finetuned on a small-scale dataset. To address this issue and integrate the OV-KWS module into different versions of the Whisper model without training, we propose an alternative to multitask finetuning, which is to use OV-KWS as an independent module and keep the Whisper model fixed. To prevent hallucination, we designed the spoken form prompt (Table 1) which incorporates disfluencies and indicated that the subsequent speech was a spontaneous talk. Compared to the naive prompt, the spoken form prompts have a more similar form to the history transcriptions used during training of the Whisper model.
3 Experimental Setups
3.1 Training configurations
For the multitask training experiments, we selected the Whisper-small model. In subsequent experiments, we also compared the Whisper-small, Whisper-medium, and Whisper-large-v2 models without finetuning, using OV-KWS as a plug-in.
For the binary CNN classifier, we employed a very lightweight architecture with four layers. Each layer had 128, 128, 256, and 256 channels respectively, and a kernel size of three. We adopted a two-phase multitask training scheduler. In the first phase, we froze the Whisper encoder and trained only the CNN classifier for OV-KWS and the Whisper decoder for ASR, for 10 epochs, with a batch size of 128 and a learning rate of $5×10^{-5}$. The total number of trainable parameters was 153.8 million (CNN 0.2 million + Decoder 153.6 million). In the second phase, we unfroze the Whisper encoder and trained the entire model for another 50 epochs, with a learning rate of $3×10^{-5}$. The number of trainable parameters increased to 241.9 million. Note that for the ASR task, we applied a random concatenation strategy [23] to increase the length of the utterance. We performed all the training on four NVIDIA T4 GPUs. During inference, we used a beam size of five for beamsearch decoding.
3.2 Datasets
To construct the training set for OV-KWS and ASR, we utilized three datasets: Aishell-1 [24], LibriSpeech [25], and TALCS [26]. Aishell-1 is a Mandarin ASR dataset with 150 hours of speech recordings. LibriSpeech is an English ASR dataset, from which we selected the train-clean subset containing 100 hours of speech recordings. TALCS is a Mandarin-English codeswitch dataset with 578 hours of speech data from English teaching scenes. For each utterance in these datasets, we randomly sampled positive words and generated speech using edge-tts  [^1] https://github.com/rany2/edge-tts. We compared the performance of the model trained with only Chinese data (zh) and the model trained with the mixture of three datasets (zh + en).
Table 2: Statistics of ASR test sets.
Dataset Utterances Duration (min) Entities
Aishell-dev 1334 113 371
Aishell-test 808 76 226
Internal-1 99 41 150
Internal-2 112 47 346
We evaluated the performance of OV-KWS and ASR on four different test sets, which consisted of two internal datasets and two open-source datasets. The internal datasets contain technical talks and manually labeled entity words that are mainly in Chinese but also include some English terms. The open-source datasets are the Aishell hot word subsets [20] in Chinese. The statistics of these datasets are shown in Table 2.
3.3 Evaluation matrics
We evaluated the performance of OV-KWS using the F1 score, which is the harmonic mean of the precision and recall. For ASR, we adopted two metrics: mixed-error-rate (MER) and entity recall. MER is a modified version of character-error-rate (CER) that can handle code-switching situations between English and Chinese. In this metric, we treat each Chinese character and each English word as a single unit. Therefore, MER is equivalent to CER for Chinese and word-error-rate (WER) for English.
4 Experimental Results
4.1 Results for open-vocabulary keyword spotting
We evaluated the performance of the CNN classifier for OV-KWS using the F1 score metric. Table 3 shows the results of different training settings. We first trained the model with a frozen Whisper encoder on the Aishell dataset only and obtained F1 scores of around 0.6 on the Aishell test sets and 0.68 on the internal datasets. Then, we evaluated the model trained with a mixture of three datasets and observed consistent improvement in the F1 scores. Finally, we unfroze the Whisper encoder, which increased the number of parameters of the OV-KWS task from 0.2 to 88.4 million and achieved significant improvement in the F1 scores. The model trained with the Aishell dataset performed slightly better on the Aishell dev and test sets, with F1 scores of 0.85 and 0.89, respectively. On the internal datasets, the model trained with the mixed dataset achieved higher F1 scores, reaching around 0.9. The comparison between single-task training and multitask training is presented in Table 4. The results indicate that jointly optimizing ASR and OV-KWS does not adversely affect the performance of the OV-KWS task and even leads to a higher F1 score on the Internal-1 dataset. Figure 2 illustrates the weight assigned to each layer of the Whisper encoder. The hidden states from the 8th to the 12th layers have a higher impact on the OV-KWS task. Moreover, the weight for the deeper layers increased slightly after finetuning, suggesting that the Whisper encoder was more adapted to the OV-KWS task.
Table 3: The performance of OV-KWS module measured by F1 score. We compared the models trained with Chinese-only data (zh) and Chinese-English-codeswitch data (zh + en). We also compared the results with the Whisper model frozen or unfrozen.
Training data zh zh + en
Unfreeze Whisper × ×
Aishell-dev 0.62 0.85 0.64 0.83
Aishell-test 0.60 0.89 0.61 0.84
Internal-1 0.68 0.75 0.71 0.90
Internal-2 0.68 0.82 0.72 0.88
Table 4: The comparison between single task training and multitask training of OV-KWS model measured by F1 score.
OV-KWS OV-KWS + ASR
Aishell-dev 0.86 0.83
Aishell-test 0.86 0.84
Internal-1 0.87 0.90
Internal-2 0.88 0.88
Figure 2: The weights for Whisper encoder layers.
Table 5: The performance of CB-Whisper measured by MER (%) / Entity Recall (%). The Whisper-small model was unfrozen and trained jointly with the OV-KWS module. "zh" means the model was finetuned on Chinese-only data. "zh+en" represents the model was trained on Chinese-English-codeswitch data. We prompted the Whisper decoder with ground-truth entities and predicted entities respectively.
Model Whisper-small (baseline) CB-Whisper-small (zh) CB-Whisper-small (zh + en)
Prompt × ground-truth × ground-truth prediction × ground-truth prediction
Aishell-dev 16.7 / 6.3 13.3 / 81.6 10.4 / 20.9 7.2 / 87.1 7.6 / 84.2 12.1 / 19.6 8.5 / 91.5 9.2 / 88.4
Aishell-test 20.1 / 10.1 15.9 / 82.1 11.6 / 14.0 8.3 / 84.4 8.6 / 82.4 13.5 / 9.8 9.7 / 91.6 10.2 / 87.7
Internal-1 5.9 / 78.5 5.3 / 95.3 14.7 / 49.0 13.5 / 76.7 13.9 / 70.2 8.9 / 61.7 8.7 / 86.8 7.7 / 86.3
Internal-2 9.7 / 60.8 8.9 / 93.7 20.0 / 32.5 17.4 / 69.5 17.2 / 64.8 16.6 / 35.9 13.0 / 76.4 13.2 / 71.4
4.2 Results for CB-Whisper
The multitask training of our CB-Whisper model resulted in the ASR performance shown in Table 5. Compared to the original Whisper-small model, the finetuned CB-Whisper-small model achieved significantly lower MER and higher Entity recall on the Aishell-dev and test sets. The model that was trained on the mixed dataset exhibited higher entity recall across all datasets, indicating better prompting function and generalisability. Remarkably, the CB-Whisper model with predicted entities surpassed the Whisper-small model with ground-truth entities in entity recall, reaching 88.4% and 87.7% on the Aishell-dev and test sets respectively. Furthermore, the CB-Whisper model outperformed the state-of-the-art SeACo-Paraformer model [20], which has entity recalls of 86% and 87% on the Aishell dev and test sets.
However, we found that though entity recall could be enhanced on internal test sets, the MERs increased slightly. This was mainly because we fine-tuned CB-Whisper on small-scale datasets, and the domain of the training data (audiobook and teaching scenes) differed from that of the test sets (technical talks). As a result, Whisper’s generalisability was reduced. Therefore, we propose a different approach, which involves freezing the original Whisper model and using OV-KWS as a separate module. As shown in Table 6, we used the Whisper-small model in two ways: either with a CNN classifier and the original whisper-small's features (the first block) or with an OV-KWS model (the second block). The second option requires more computation, as two different whisper-encoders were used, but it can achieve higher entity recalls. A benefit of this method is that the OV-KWS model can be applied to various Whisper models without fine-tuning. However, we found that with the naive prompt, the MER can increase compared to the original model. The main reason was that the Whisper model tended to omit filler words and disfluencies if the prompt was in a well-formatted text form. This problem could be alleviated by using spoken form prompts that mimicked a presentation transcription. Compared to the naive prompt and no prompt, spoken form prompts consistently improved the MER. As an example, using the OV-KWS model and the spoken form prompt, the MER decreased from 4.4% to 2.9% and from 5.6% to 3.6% on the Internal-1 and Internal-2 subsets respectively for the Whisper-large-v2 model  [^2] We have conducted experiments with the Whisper-large-v3 model, but it did not outperform the v2 model.. In addition, the proposed method attained entity recalls of 97.7% and 96.9% for the two datasets, which are considerably superior to the original model’s performance of 87.3% and 79.9%, respectively.
Table 6: The performance of various Whisper models with OV-KWS as an independent module measured by MER (%) / Entity Recall (%). We compared the naive prompt and the spoken form prompt.
Model Prompt Internal-1 Internal-2
CB-Whisper-small × 5.9 / 78.5 9.7 / 60.8
(frozen) Naive 5.5 / 91.5 8.5 / 87.9
Spoken 4.2 / 91.2 6.6 / 89.0
Whisper-small Naive 5.4 / 94.0 9.0 / 90.2
+ OV-KWS Spoken 4.0 / 95.9 6.8 / 91.1
Whisper-medium × 4.8 / 84.2 6.7 / 70.4
+ OV-KWS Naive 6.8 / 94.3 14.6 / 90.3
Spoken 4.0 / 95.6 4.8 / 94.4
Whisper-large-v2 × 4.4 / 87.3 5.6 / 79.9
+ OV-KWS Naive 4.7 / 97.4 6.6 / 95.1
Spoken 2.9 / 97.7 3.6 / 96.9
5 Conclusion
This paper presents CB-Whisper, an innovative ASR framework that leverages the prior knowledge of entity words to enhance their recalls. Our model consists of an OV-KWS module that exploits the hidden states of the Whisper encoder. The entities detected by this module serve as prompts for the Whisper decoder. We devise a multitask training strategy that simultaneously optimizes OV-KWS and ASR tasks. We also propose a more convenient method to apply the OV-KWS model to various Whisper models without finetuning using spoken form prompts. We evaluate our method on four test sets and show that it achieves substantial improvements in both MER and entity recall. In future work, we aim to enhance the efficiency of the OV-KWS module to enable it to handle a large keyword database and conduct large-scale experiments by distilling speech data from the original Whisper model, to boost the performance of Multitask finetuning in the general domain.